2+1 REMPI OF CO AND N_{2} OBSERVED OPTOGALVANICALLY IN GLOW DISCHARGES

\underline{IAN MACGREGOR}, STILIANA ANTONOVA, ANTONIS P. TSAKOTELLIS, AO LIN, AND GEORGE C. MCBANE, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

2+1 resonance enhanced multiphoton ionization spectra of CO (through the B^{1}Σ^{+} intermediate state) and N_{2} (through the a′′ ^{1}Σ_{g}^{+} state) have been observed optogalvanically. UV light pulses focused into the positive columns of 1–10 Torr glow discharges produced easily observable changes in the discharge currents. The signal intensities appeared to provide quantitative measurements of the vibrational and rotational population distributions in the discharges. Population analyses and comparisons with independent population measurements will be presented.